OpenStudy (anonymous):

Simplify (This is rational equations) (2x^2 + 8x + 4) / (x (x + 1)(x + 2)) Simplify (This is rational equations) (2x^2 + 8x + 4) / (x (x + 1)(x + 2)) @Mathematics

OpenStudy (asnaseer):

\[\frac{2x^2+8x+4}{x(x+1)(x+2)}=\frac{2(x^2+4x+2)}{x(x+1)(x+2)}\]

OpenStudy (anonymous):

Do I also have to factor the polynomial on the top numerator?

OpenStudy (anonymous):

Lol my bad you cant >.<

OpenStudy (asnaseer):

into cannot be factored as such. the only other simplification I can think of is:\[\frac{2(x^2+4x+2)}{x(x+1)(x+2)}=\frac{2((x+2)^2-2)}{x(x+1)(x+2)}=\frac{2(x+2)}{x(x+1)}-\frac{4}{x(x+1)(x+2)}\]

OpenStudy (anonymous):

Oh ok thank you so much

OpenStudy (asnaseer):

no problem